# UnZipLoRA: Separating Content and Style from a Single Image

### v1

@@ 4.4 Limitations
Figure 8: Failure Cases. For a few input images with abstract styles, UnZipLoRA fails to accurately destylize the subjects. Note the artifact introduced on kangaroo’s leg due to curved line-art style, and unnatural shape of the bear due to high levels of abstraction.
5 Conclusion and Future Work
@@
We introduced UnZipLoRA, a novel method for decomposing a single image into its subject and style, represented as two disentangled and compatible LoRAs. By employing prompt, column, and block separation strategies, UnZipLoRA effectively extracts and preserves these elements, enabling diverse recontextualizations and manipulations. Our experiments demonstrate superior performance compared to existing methods, highlighting UnZipLoRA’s potential for creative exploration and control within text-to-image generation. Future work includes exploring alternative disentanglement techniques and training-free approaches for subject-style decomposition.
@@ A Additional Implementation Details
Block separation strategy. As discussed in Sec. 3.3.3, we employ block separation technique similar to that proposed in B-LoRA. Specifically, as shown in Fig. 9, the U-Net architecture in SDXL comprises three primary components: the downsampling blocks, the middle blocks, and the upsampling blocks. Each of these components contains several groups of transformer-based blocks. As the upsampling component is more critical to the overall performance, our primary focus for block separation lies in upsampling. Within the upsampling component, there are two distinct groups of blocks, which are differentiated by number of transformers per block: the first group (Upblock0 in Fig. 9) contains blocks with 10 transformers each, while the second group (Upblock1) has blocks with only 2 transformers. Due to this disparity, the first group plays a more significant role in the learning process.
@@
Training prompts. As mentioned in the main paper, we use three separate prompts during training: a combined prompt, and two separate prompts (for subject and style each). Our combined prompt is of the form “a <$c$> in <$s$> style”, while “A <$c$> ” and “An image in <$s$> style” are used as subject-specific and style-specific prompts respectively. Here, <$c$> and <$s$> represents the subject and style descriptors specific to the input image. We follow the descriptor strategies of existing approaches such as ZipLoRA [32] and StyleDrop [33] that advocates using a unique token (such as “sks”) to describe the subject, and use detailed artistic description for the style. For example, the prompt we use for the first input image in Fig. 12 is “A sks sun hat in flat cartoon illustration style”. As discussed in DreamBooth [29], using an unique token helps the model to bind the new subject to a unique embedding vector in space since a subject is usually spatially localized in a particular region of the image. On the other hand, style is typically spread throughout the image, thus using detailed text description increases the flexibility during style training as reported by StyleDrop [33]. To ensure a fair comparison between our method and others, such as B-LoRA, we use the same prompt formations for all the competing methods.
Figure 10: Ablation Examples. We show the impact of adding prompt-wise separation, column-wise separation, and block-wise separation sequentially. Each row illustrates the outputs of baseline methods and our proposed approaches, highlighting their difference in subject-style disentanglement, style fidelity, and recontextualization across different examples.
B Compute Requirements
C Additional Results
C.1 Ablation Examples
We provide additional examples depicting the effects of adding each component of our method, as shown in Fig. 10. Each row demonstrates the recontextualization of subject, style and subject-style across different configurations and baseline methods.
Figure 11: The interface used for user studies. We depict the interface we used for method comparison user study shown in the main paper.
@@ C.1 Ablation Examples
C.2 Subject-Style Recontextualization Comparison
We present additional examples of subject-style recontextualization using diverse prompts and input images comparing our method with B-LoRA in Fig. 12. The results further demonstrate our method’s superiority in preventing overfitting, reproducing accurate subject and style representations, and enabling flexible recontextualization.
Prevention of overfitting. Our method mitigates overfitting by disentangling subject and style representations, ensuring diverse and robust outputs even with challenging prompts
@@ C.2 Subject-Style Recontextualization Comparison
C.3 Additional Qualitative Comparisons
@@
B-LoRA. As what we discussed in Experiments section of main paper, and in Sec. C.2, B-LoRA fails to generate consistent results and suffers from overfitting to input images. While it can always accurately learn style features, it struggles to reproduce subject details reliably. For instance, in the guitar and sunflower examples, B-LoRA fails to consistently retain the original input’s color.
@@ D Image Attributions
D.1 Image attributions for the stylized images used as inputs

### v2

@@ 4.5 Applications and extensions
Figure 8: Failure Cases. In a few cases with highly abstract styles, UnZipLoRA may fail to destylize the subject accurately (note the unnatural shape of the bear). B-LoRA exhibits a similar failure.
5 Conclusion
@@
This paper introduced UnZipLoRA, a novel method for decomposing a single image into disentangled, compatible subject and style LoRAs. UnZipLoRA utilizes prompt, column, and block separation strategies to effectively extract these elements, enabling diverse recontextualizations and manipulations. Our experiments demonstrate superior performance compared to existing methods, highlighting UnZipLoRA’s potential for creative exploration and control within text-to-image generation. While robust across a wide range of subjects and styles, UnZipLoRA may fail to destylize the subject when the style involves too much shape distortion or abstraction, as observed in Fig. 8. Future work includes exploring alternative disentanglement techniques for such challenging cases, training-free approaches for improved efficiency, and extending UnZipLoRA to other architectures for better generalization.
Acknowledgments. This research was supported in part by NSF grant CCF 2348624.
@@ A Additional Implementation Details
Block separation strategy. As discussed in Sec. 3.3.3, we employ a block separation technique similar to that proposed in B-LoRA. Specifically, as shown in Fig. 9, the U-Net architecture in SDXL comprises three primary components: the downsampling blocks, the middle blocks, and the upsampling blocks. Each of these components contains several groups of transformer-based blocks. As the upsampling component is more critical to the overall performance, our primary focus for block separation lies in upsampling. Within the upsampling component, there are two distinct groups of blocks, which are differentiated by number of transformers per block: the first group (Upblock0 in Fig. 9) contains blocks with 10 transformers each, while the second group (Upblock1) has blocks with only 2 transformers. Due to this disparity, the first group plays a more significant role in the learning process.
@@
B Additional Experiments and Results
B.1 Trigger Phrases Selection
As mentioned in the main paper, we follow the standard prompt construction strategy “a <cc> in <ss> style” with trigger phrases <cc> and <ss> in our text prompt to describe the content and style respectively. For selecting the trigger phrases <cc> and <ss> , we follow the descriptor strategies of existing approaches such as DreamBooth [30], StyleDrop [34], ZipLoRA [33], and B-LoRA [6].
Figure 10: Trigger phrase selection for the subject.
Figure 11: Trigger phrase selection for the style.
Figure 12: User Study Interfaces. We depict the graphical user interfaces (GUI) we used in (left) methods comparison user study for comparing quality of subject-style decomposition, and (right) ablations user study for comparing subject decomposition, style decomposition, and combined recontextualization. User study results are included in the main paper.
@@ B.1 Trigger Phrases Selection
As discussed in DreamBooth [30], using a unique token for the subject helps the model to bind the new subject to a unique embedding vector in space since a subject is usually spatially localized in a particular region of the image. On the other hand, style is typically spread throughout the image, thus using just a generic text description increases the flexibility during style training as reported by StyleDrop [34]. This choices are further validated by the follow up works such as ZipLoRA [33] and B-LoRA [6]. Thus, in our experiments we use a unique token followed by the subject class label as the subject trigger phrase (e.g. <cc> =‘sks dog’), and use generic artistic description for the style (e.g. <ss> = ‘watercolor painting’). For example, the prompt we use for the first input image in Fig. 13 is “A sks sun hat in flat cartoon illustration style”.
Here we independently validate the above choices by studying the impact of the trigger phrases on the results of our method as the degree of detail of these phrases is varied. We confirm that a single-word class label for the subject, and a generic, brief (2-3 word) description for style are sufficient to effectively guide our method, while a more detailed prompt provides additional flexibility for reinforcing any desired attributes.
In Fig. 10 and Fig. 11, we show several groups of examples demonstrating the influence of subject and style trigger phrases on generation. We compare results trained with more general prompts, such as referring to a subject by its category rather than its specific name or using broad and vague style phrases (e.g. using ‘flower’ instead of ‘sunflower’). The results indicate that the generation quality with general prompts is largely preserved, showing no noticeable degradation compared to more detailed prompts.
In some cases, using detailed phrases can help reinforce specific attributes, thus providing flexibility to users. For example, in Fig. 10, using ‘flower’ as a trigger phrase retains most of the characteristics of the input subject, while using a more detailed prompt ‘sunflower’ helps boost the fidelity of the features in the center of the flower. Similarly, using ‘microscope’ instead of generic ‘machine’ helps retain the fine-grained shape and color characteristics of its eyepiece tube.
Similar conclusions hold for style trigger phrases as well (see Fig. 11): UnZipLoRA can successfully learn the style of the input image even with highly generic, single word style phrases such as ‘cartoon’ and ‘painting’. At the same time, more detailed descriptions can help reinforce specific attributes. While these attributes may still be captured without explicit mention, incorporating them into the prompt ensures more stable preservation. For example, the differences in generations of ‘golden’ style in Fig. 11 are subtle, yet a more artistically descriptive phrase ‘melting golden 3d rendering’ leads to clearer stylistic features such as flowing golden droplets. Another example is the pizza in Fig. 11: UnZipLoRA works well with a generic phrase ‘cartoon’, and specifying ‘ukiyo-e’ helps retaining more of the background waves, and produces a stronger ‘ukiyo-e’ aesthetics.
These findings suggest that general trigger phrases suffice for capturing the overall subject and stylistic impression. When retaining specific features is required, explicitly incorporating those features into the input prompt is beneficial.
B.2 Subject-Style Recontextualization Comparison
A key advantage of UnZipLoRA is its ability to produce compatible subject and style LoRAs that can be seamlessly merged via direct addition. This allows for the generation of novel recontextualizations that faithfully incorporate both the subject and style of the original image. Figure 3 in the main paper indemonstrates this capability through various recontextualizations using either individual LoRAs or a combination of both. In this section, we provide qualitative comparisons between our method and B-LoRA for the task of subject-style recontextualization. As demonstrated in Fig. 13, UnZipLoRA is superior in preventing overfitting, reproducing accurate subject and style representations, and enabling flexible recontextualization.
Preventing overfitting. Our method mitigates overfitting by disentangling subject and style representations, ensuring diverse and robust outputs even with challenging prompts
@@ B.2 Subject-Style Recontextualization Comparison
Figure 15: Additional Decomposition and Re-contextualization Results. We present additional results on diverse inputs such as humans and 3D rendering styles. The results showcase UnZipLoRA’s ability to preserve the details of both the subject and style along with providing flexible recontextualizations.
B.3 Additional Qualitative Comparisons
@@
B-LoRA. As what we discussed in Experiments section of main paper, and in Sec. B.2, B-LoRA fails to generate consistent results and suffers from overfitting to input images. While it can always accurately learn style features, it struggles to reproduce subject details reliably. For instance, in the guitar and sunflower examples, B-LoRA fails to consistently retain the original input’s color.
@@
B.4 Additional Qualitative Results on Diverse Inputs
We present additional decomposition and recontextualization results using a wider variety of subjects/styles, including more complex subjects such as humans and styles such as 3D rendering in Fig. 15. The results demonstrate that our method generalizes effectively to detailed and diverse subjects, maintaining the fidelity of the subjects/styles, and flexibly recontextualizing them across various contexts, showcasing the superiority of our method in handling rich and challenging inputs.
B.5 Additional Ablation Study Results
Figure 16: Ablation study. We show the impact of adding prompt-wise separation, column-wise separation, and block-wise separation sequentially. Each row illustrates the outputs of baseline methods and our proposed approaches, highlighting their difference in subject-style disentanglement, style fidelity, and recontextualization across different examples.
Figure 17: Additional Results on KOALA Diffusion. Our approach generalizes effectively, as demonstrated by successful results on the more recent KOALA diffusion model.
We provide additional examples of ablation study of the effects of adding each component, as shown in Fig. 16. Each row demonstrates the recontextualization of subject, style and subject-style across different configurations and baseline methods.
Table 4: Ablation Study Alignment Scores. Comparisons for Content and Style Decomposition among each separation strategy.
DB-LoRA M1 M2 M3 (UnZipLoRA)
Style-align. (CLIP-I) $\uparrow$ $0.417$ $0.409$ $0.407$ $\mathbf{0.427}$
Subject-align. (DINO) $\uparrow$ $0.339$ $0.346$ $0.347$ $\mathbf{0.349}$
Style-align. (CSD) $\uparrow$ $0.245$ $0.217$ $0.216$ $\mathbf{0.265}$
Subject-align. (CSD) $\uparrow$ $0.338$ $0.352$ $0.354$ $\mathbf{0.358}$
Quantitative results. Beyond the ablation user study results provided in Tab. 3 in the main paper, we provide additional quantitative results in the form of subject- and style-alignment scores in Tab. 4. The results confirm each separation strategy’s independent contribution: prompt separation (M1) aids subject learning, column separation (M2) improves disentanglement, and block separation (M3) significantly enhances stylization. These trends align with the user study results (Tab. 3) and our analysis in Sec. 4.4 of the main paper.
User study interface. We conducted user studies in the main paper to compare our method with the competing approaches. Beyond the configurations, results, and analyses presented in the main paper, we include the interface used for the main user study and the ablation user study in Fig. 12.
Figure 18: Ablations on different values of $N$. Our column selection strategy selects top $N\%$ of most important columns usiing dynamic importance recalibration strategy. As shown, $20\leq N\leq 40$ is sufficient for capturing subject/style successfully.
B.6 Ablation on Percentage of High-importance Columns
To evaluate the influence of the column selection percentage (N) within our dynamic importance recalibration strategy, we conducted ablation studies across varying values of $N$. This strategy selects the top $N\%$ of most important columns by calculating the column importance from the gradient information using the Cone method [22]. The results of our ablation experiments shown in Fig. 18 confirm that a selection percentage within the range of $20\leq N\leq 40$ is sufficient for successfully capturing both subject and style characteristics.
B.7 Additional Cross-combination Results
The subject and style LoRAs produced by UnZipLoRA open up a possibility for cross-combination: pairing a subject LoRA from one image with a style LoRA from another. Fig. 19 provides additional results for such cross-combination where the LoRAs are combined by direct addition. While these LoRAs are not explicitly trained together (and thus not subject to the orthogonality constraints enforced by ZipLoRA [33]), the inherent separation imposed by our column and block strategies generally results in higher compatibility than generic DreamBooth-LoRAs trained without such constraints. Consequently, direct arithmetic merger yields promising cross-stylization results.
Figure 19: Additional examples of cross-composition using subject and style from different input images. The result demonstrates that our method effectively integrates features from both subject and style.
B.8 Additional Results on KOALA Diffusion
Our method extends beyond SDXL, and is applicable to newer diffusion models. We demonstrate this generalizability by training on KOALA [19], a more efficient, recent text-to-image model with leaner architecture compared to SDXL. We provide additional results in Fig. 17. As shown, our method, when applied to KOALA, accurately captures subject and style and allows for successful recontextualization (though the overall quality of the results is not as high as for SDXL due to limited capacity and lower parameter count of KOALA).
C Compute Requirements
@@ D Image Attributions
Image attributions for the stylized images used as inputs